1. Valles de Benasque y Eriste

BENÀS (BENASQUE), ANSILS (ANSILES), SARLLÉ (CERLER), BENÀS 1 h. 45 minutes (one way) 2 h. 35 min. (complete excursion)

Leave Benàs by the Ansils road. You will get to Ansils in about 20 minutes. Once in the town of Ansils you will see the old track to Sarllé, indicated by green arrows. Take it. After some minutes of walking you have to pass by the graveyard. You will find four forks on your way up; always take the one on the right. After 1 hour and 15 minutes you will cross a trail. Take it to the right and in about 10 minutes you will get to the lifts of the ski trails of Cerler. You can go back to Benasque by the old way to Cerler (consult the previous page).